the perimeter of a triangle is 73 inches. If two sides are of equal length and the third side is 7 inches longer than the others, find the length of the three sides
Found 2 solutions by fractalier, addingup:
Answer by fractalier(6550)

Call one of the equal sides x. Then the three sides are
x,x, and x+7...they add to 73, so
X + x + x + 7 = 73
3x + 7 = 73
3x = 66
x = 22
The sides are 22, 22, and 29.

Answer by addingup(3677)

Let's call the unknown sides x. Since we have two sides, we have 2x.
And the long side, the problem says, is x+7
2x+x+7= 73 Add x on left, and subtract 7 on both sides
3x= 66 Divide both sides by 3
x= 22 is the length of each short leg, and the long leg is
22+7= 29
Proof
2 equal sides:
2*22= 44 Plus the long side:
22+7= 29
_ _ -------
Total: 73 We have the correct answer